OpenStudy (anonymous):

Graph the line passing through the given point and having the given slope. Give the slope intercept form of the equation of the line if possible. (-1, 3), m= -3/2 Post@Rachella

OpenStudy (anonymous):

I want to know how to I find the points

OpenStudy (anonymous):

You're wanting to know how to find the points on the graph?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know how to setup a problem like this

OpenStudy (anonymous):

Are u there

OpenStudy (anonymous):

Given two points: \[(x_{1},y _{1})~and~(x _{2,}y _{2})\]The formula for the slope of the straight line going through two points are: \[m = \frac{ y _{1}-y _{2} }{ x _{1}-x _{2} }\]Where the subscripts merely indicate that you have a "first" point (whose coordinates are subscripted with a "1") and a "second" point (whose coordinates are subscripted with a "2"); that is, the subscripts indicate nothing more than the fact that you have two points to work with. Note that the point you pick as the "first" one is irrelevant; if you pick the other point to be "first", then you get the same value for the slope:\[m = \frac{ y _{2}-y _{1} }{ x _{2} -x _{1}}\](If you're not sure that the two formulas above give exactly the same values, no matter the pair of points plugged into them, then pick some points and try them out. See what you get.) The formula for slope is sometimes referred to as "rise over run", because the fraction consists of the "rise" (the change in y, going up or down) divided by the "run" (the change in x, going from left to the right). If you've ever done roofing, built a staircase, graded landscaping, or installed gutters or outflow piping, you've probably encountered this "rise over run" concept. The point is that slope tells you how much y is changing for every so much that x is changing.

OpenStudy (anonymous):

Yeah I know this formula

OpenStudy (anonymous):

Which formula are you wanting to know then?

OpenStudy (anonymous):

I want to know how to solve this problem I wrote

OpenStudy (anonymous):

Are you having problems figuring out how to solve to this

OpenStudy (anonymous):

Alright. Let's look at your points. The equation of line passing through point: \[A = (-1,3) ~and ~point~B = (-3,2)\]Is: \[y = \frac{ 1 }{ 2 }x+\frac{ 7 }{ 2 }\]To find the equation of the line passing through points: \[A(x _{A},y _{A})~and~B(x _{B},y _{B})~(x _{A}\neq~x _{B})\]We use the formula provided in my previous comment. In this example we have: \[x _{A}= -1, ~y _{A}=-3,~and~y _{B}=2\]So with this being said, \[y - y _{A}= y _{B}-y _{A}~\over~x _{B}-x ^{}\]\[y - 3 = \frac{ 1 }{ 2 }(x+1)\]\[y -3 = \frac{ 1 }{ 2 }x+\frac{ 1 }{ 2 }\]Then to get your final answer: \[y = \frac{ 1 }{ 2 }x + \frac{ 7 }{ 2 }\]

OpenStudy (anonymous):

Sorry that it took so long! Typing all of those equations then getting OpenStudy notifications saying "Unable to connect ... trying to reconnect" makes it a process. There are the steps provided in order, the way I got the answer. I hope you understand and this helps! Again, sorry that it took so long!